Tohban Report 2012-12-12

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 05 Dec 2012
End Date: 12 Dec 2012
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

This week was very quiet, with a smooth B-level background for most of the week. There were a few C-class events that are well observed relative to the quiet background. There are currently three non-complex sun spots near disk center and two that are rotating off. Stereo B shows one spot that will rotate on in the coming week.

How many GOES flares occurred?

 Flares above B, C, M, X class were     39     4     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     23     3     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     13     1     0     0

There were RHESSI flares/GOES flares 85 / 43 over the time range 04-Dec-12 11-Dec-12

Memory Management

This week's SSR shot up to ~70% and would empty to ~40% over the course of three days, while the cause of the issue was being determined. The drastic rise in SSR was coincident with the commanding to start taking rear data through the entire night, but it was determined that these two events were not related.

D8's decay out of the SAA has gotten progressively worse in the past week, occasionally taking the entire next pass to decay to counts recorded by the other detectors. D8 events were turned off and at the first pass this morning the SRR emptied by 30%, indicating that it is the source of the memory issue. Please see the Detector Issues section for more information.

Spacecraft Status

Brian Dennis showed a plot which indicates that the temperature rise is decreasing. Earlier in the year the temperature was increasing by 10k/9months and now we've had a 2K/4month increase since july. Before the anneal the temperature was rising by .33K/month.

Data Gaps

There are two data gaps in VC3 data that cannot be replayed due to the SSR fill issue.

GAP START TIME GAP END TIME GAP (SEC) 2012-12-09T05:50:00.000 -- 2012-12-09T05:55:00.000 300.00000

GAP START TIME GAP END TIME GAP (SEC) 2012-12-07T09:55:00.000 -- 2012-12-07T10:10:00.000 900.00000

Detector issues

To mitigate the D8 SSR issue, events have temporarily been turned off to empty the SSR. D8's voltage was lowered by 10 steps on Monday Dec 12th at 21:39. This has not alleviated the problem and the voltage will be lowered by another 10 steps this week.

D8 is currently sitting at 3000V, so there is room to try lowering the voltage to fix the long decay out of SAA. Since more radiation damage occurs at lower bias voltages, there is concern about lowering the voltage too far. Lab experiments suggest that radiation damage at lower voltages will persist even when the voltage is increased. In practice, RHESSI hasn't seen evidence of this. David suggested slowly lowering the voltage by 10 steps in a pass and monitoring the state of the detector.

David also mentioned the possibility of raising D8 voltage or raising its decimation to fix the counts problem.

This week it was discovered that D4 has been spontaneously segmenting and desegmenting for approximately the past 30 days. To stabilize its segmentation, its HV was raised by 5 steps. It appears as though D4 segmentation is now stable, but tohbans should still monitor its state.

Brian Dennis' D4 powerpoint at the end of the RHESSI meeting indicated that the current calibration of D4 is off, and most likely its referenced to the unsegmented D4 calibration. He also showed that its resolution is as good as it was before the anneal.

Other notes

Decimation was changed from Normal/Vigorous to Active/Vigorous in an effort to combat the SSR fill problem. It remains in Active/Vigorous at this time.

Night time data was turned off completely.

Added a page to the wiki showing the ADC table for the nine detectors: "RHESSI Detector Channel vs Energy"

Spacecraft Management

Decimation active/vigorous
HLAT Decimation
Night time data (fronts) turned off
Night time data (rears) turned off
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? See above.
Personal tools